Wrongly called for a foul, an angry basketball player throws the ball straight down to the floor. If the ball bounces straight up and returns to the floor 2.7 after first striking it, what was the ball's greatest height above the floor?
Answer by Alan3354(69443)

Wrongly called for a foul, an angry basketball player throws the ball straight down to the floor. If the ball bounces straight up and returns to the floor 2.7 after first striking it, what was the ball's greatest height above the floor?
--------------------
h(t) = -16t^2 + vt (v is the upward speed as it leaves the floor)
units are feet, using 32 feet/sec/sec for gravity.
--------------------
To find v, the initial velocity, set h(t) = 0 and t = 2.7 seconds.
0 = -16*7.29 + v*2.7
v = 16*2.7 = 43.2 ft/sec
------------------------
Since the ball takes the same amount to time to rise as to descend, it is at its max height at 1/2 of the 2.7 seconds, = 1.35 seconds
h(1.35) = -16*1.35^2 + 43.2*1.35
h = 29.16 feet max height